OpenStudy (anonymous):

Consider the equation x_1 + x_2 + x_3 + x_4 = n where X_i >= 0 integers. x_2 >=2 x_3 is multiple of four and 1<=x_4<=3. Let Cn be number of solutions to this equation subject to these restrictions. find generating function for the sequecne {c_n : n>=0} and use it to find closed formula for c_n

OpenStudy (perl):

what is a generating function (what class is this )

OpenStudy (anonymous):

Lol combinatorics... I am assuming this comment is a joke ;)

OpenStudy (anonymous):

X_1 is 1/1-x X_2=x^2/1-x X_3= 1/1-x^4 X_4= (x(1-x^3))/1-x I am figuring out where I go next..